### v2

# Statistical Test on Diffusion Model-based Generated Images by Selective Inference

Abstract
AI technology for generating images, such as diffusion models, has advanced rapidly. However, there is no established framework for quantifying the reliability of AI-generated images, which hinders their use in critical decision-making tasks, such as medical image diagnosis. In this study, we propose a method to quantify the reliability of decision-making tasks that rely on images produced by diffusion models within a statistical testing framework. The core concept of our statistical test involves using a selective inference framework, in which the statistical test is conducted under the condition that the images are produced by a trained diffusion model. As a case study, we study a diffusion model-based anomaly detection task for medical images. With our approach, the statistical significance of medical image diagnostic outcomes can be quantified in terms of a $p$-value, enabling decision-making with a controlled error rate. We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
1 Introduction
Advances in image generation AI, such as diffusion models, have been remarkable (Song and Ermon, 2019). They can generate highly realistic and detailed images, which leads to innovations in various tasks across various fields. For example, image generation AI significantly enhances medical image diagnosis by improving accuracy and efficiency. It can generate highly detailed and enhanced images from standard medical scan images, potentially offering doctors to detect anomalies and diseases with greater precision. Furthermore, image generation AI can be used to create alternative versions of medical images to consider what-if scenarios. For example, it can generate virtual images of a patient when they are healthy, which allows for comparing the current actual images with the virtual healthy images, making it possible to provide a diagnosis tailored to the individual patient.
On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions. Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, it may inherently contain biases and errors. Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes. When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI. However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test. When a test image, which may contain an abnormal region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process. By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated. By comparing the input image with the normal image, the anomaly region can be identified. In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomaly regions in the form of $p$-value. The proposed DMAD-test calculates the $p$-values by incorporating the fact that the anomaly region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
In this study, we address this challenge using the statistical hypothesis testing framework. We introduce a statistical inference framework called selective inference (SI), which has gained attention over the past decade in the statistics community as a novel approach for data-driven hypotheses (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014). In SI, statistical inference is performed based on the sampling distribution of the test statistic under the condition that the hypothesis being tested was selected based on the data. Our core idea is to formulate decision-making tasks involving generated images as statistical hypothesis testing problems, and to incorporate SI framework to accurately quantify the reliability of decisions influenced by these generated images.
As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomaly regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1). Initially, a diffusion model is trained exclusively on normal images during the training phase. In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomaly regions. Our proposed statistical test, the Diffusion Model-based Anomaly Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values. Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05) even in finite sample settings.
Related works.
Diffusion models have been effectively utilized in anomaly region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains abnormal regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomaly regions. Other types of generative AI has also been used for anomaly region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them work on image generation by diffusion models.
Contributions.
Our main contributions in this study are summarized as follows. Our first contribution is the introduction of a statistical testing framework for quantifying reliability in decision-making based on images generated by diffusion models. The second contribution is the implementation of SI for diffusion models, which requires the calculation of the sampling distribution conditional on the diffusion model, necessitating the development of non-trivial computational methodology. The third contribution is to theoretically guarantee the performance of the proposed DMAD-test and demonstrate its performance through numerical experiments and applications in brain imaging diagnostics. The code is available as supplementary material.
2 Diffusion Models
In this section, we briefly explain the diffusion model employed in this study. Given a test image which possibly contain abnormal regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{(t\geq 1)}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$
where
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\bm{I}).$
By the reproducibility of the Gaussian distribution, $\mathbf{x}_{t}$ can be rewritten by a linear combination of $\mathbf{x}_{0}$ and $\epsilon$, i.e.,
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I}),$ (1)
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}\bm{I})$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon||^{2}_{2}$. Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t},$ (2)
where
$f^{(t)}_{\theta}(\mathbf{x}_{t}):=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}},$ (3)
and
$\sigma_{t}=\eta\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}.$ (4)
Here, $\eta$ is a hyperparameter that controls the randomness in the reverse process. By setting $\eta=1$, we can create new images by stochastic sampling. On the other hand, if we set $\eta=0$, deterministic sampling is used for image generation. By recursively sampling as in (2), we can obtain a reconstructed image of the original input $\mathbf{x}_{0}$.
In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the abnormal regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the abnormal region to the normal region. Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost. For the purpose of reducing computational cost, various methods have been proposed. For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A). The image reconstruction scheme by DDPM is summarized in Algorithm 1.
Algorithm 1 Reconstruction Process
0:  Input image $\mathbf{x}$
1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon_{T^{\prime}}$
2:  for $t=T^{\prime},\dots,1$ do
3:     $f^{(t)}_{\theta}(\mathbf{x}_{t})\leftarrow(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$
4:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})+\sigma_{t}\epsilon_{t}$
5:  end for
5:  Reconstructed image $\mathbf{x}_{0}$
3 Statistical Test on Generated Images by Diffusion Models
In this section, we formulate the statistical test for detecting anomaly regions using images generated by a trained DDPM model. As shown in Figure 1, anomaly region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomaly regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomaly regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomaly regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomaly regions can be detected by comparing the original test image and its reconstructed one.
Problem Formulation.
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret that an image is considered as a sum of true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and noise component $\bm{\epsilon}\in\mathbb{R}^{n}$. Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\epsilon},\;\;\;\bm{\epsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix, which is known or estimable from external data. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}^{\rm obs}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomaly regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
$E:\mathbb{R}^{n}\ni\bm{X}\mapsto|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|\in\mathbb{R}^{n},$ (6)
where absolute value is taken pixel-wise. Anomaly regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomaly region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
$\mathcal{M}_{\bm{X}}=\left\{i\in[n]\mid E_{i}(\bm{X})\geq\lambda\right\}.$ (7)
Statistical Inference.
In order to quantify the statistical significance of the anomaly regions detected by a diffusion model, let us consider a statistical test with the following null and alternative hypotheses:
$\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i}~{}~{}~{}\text{v.s.}~{}~{}~{}\mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i},$ (8)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the anomaly regions and other regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (8) is the difference in mean pixel values between the inside and outside of the anomaly region, i.e.,
$T(\bm{X})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\sum_{i\in\mathcal{M}^{c}_{\bm{X}}}X_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\bm{X},$ (9)
where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}^{c}_{\bm{X}}}$ is the vector that depends on the anomaly region $\mathcal{M}_{\bm{X}}$ and $\mathbf{1}^{n}_{\mathcal{C}}\in\mathcal{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
If we do not account for the fact that the anomaly region is detected by a diffusion model, the distribution of the test statistic would be simply given as $T(\bm{X})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. In this case, the $p$-values defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\right),$ (10)
would be easily computed by the normality of the test statistic distribution. However, in reality, since the anomaly region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated. Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.
4 Computing Selective $p$-values
In this section, we introduce selective inference (SI) framework for testing images generated by diffusion models and propose a method to perform valid hypothesis test.
4.1 Conditional Distribution of Test Statistics
Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X})$. Then, we consider the sampling distribution of $T(\bm{X})$ conditional on the event that the anomaly region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomaly region $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$, i.e.,
$T(\bm{X})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (11)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameters that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The nuisance parameter $\mathcal{Q}_{\bm{X}}$ is written as
$\mathcal{Q}_{\bm{X}}=\left(I_{n}-\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\bm{X}.$ (12)
The $p$-value calculated from this conditional sampling distribution is called a selective $p$-value. Specifically, the selective $p$-value is defined as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (13)
where $\mathcal{X}$ is the conditional data space defined as
$\mathcal{X}=\left\{\bm{X}\in\mathbb{R}^{n}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (14)
Due to the conditioning on the nuisance parameter $\mathcal{Q}_{\bm{X}}$, the conditional data space $\mathcal{X}$ can be rewritten as
$\mathcal{X}=\left\{\bm{X}(z)\in\mathbb{R}^{n}\mid\bm{X}(z)=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (15)
where vectors $\bm{a},\bm{b}\in\mathbb{R}^{n}$ are defined as
$\bm{a}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}},\;\bm{b}=\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}},$ (16)
and the region $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (17)
Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z^{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}+\bm{b}Z$ and $\bm{X}^{\mathrm{obs}}=\bm{a}+\bm{b}z^{\mathrm{obs}}$. Then, the selective $p$-value in (13) is re-written as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (18)
Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution. Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (18) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
4.2 Over-Conditioning
To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach. It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model. The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). It is easy to understood that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$. Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomaly region) as the observed one is selected. In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$. By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
Identification of $\mathcal{Z}^{\mathrm{oc}}$.
Let us write a polyhedron $\mathcal{P}$ composed of piecewise-linear functions as
$\mathcal{P}_{k}=\left\{\bm{\Delta}_{k}\bm{X}\leq\bm{\delta}_{k}\right\},k\in[K],$ (19)
where $\bm{\Delta}_{k}$ and $\bm{\delta}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions of the $k$-th piecewise-linear function, respectively. Then, a piecewise-linear function $\mathcal{A}(\bm{X})$ is written in the following form:
$\mathcal{A}(\bm{X})=\begin{cases}\bm{\Psi}_{1}\bm{X}+\bm{\psi}_{1}&\text{if }\bm{X}\in\mathcal{P}_{1},\\ \bm{\Psi}_{2}\bm{X}+\bm{\psi}_{2}&\text{if }\bm{X}\in\mathcal{P}_{2},\\ \hskip 28.45274pt\vdots\\ \bm{\Psi}_{K}\bm{X}+\bm{\psi}_{K}&\text{if }\bm{X}\in\mathcal{P}_{K},\end{cases}$ (20)
where $\bm{\Psi}_{k}$ and $\bm{\psi}_{k}$ for $k\in[K]$ are the coefficient matrix and the constant vector with appropriate dimensions for the $k$-th polyhedron, respectively. Using the notation in (15), since the input image $\bm{X}(z)$ is restricted on a one-dimensional line, each component of the output of $\mathcal{A}$ is written as
$\begin{gathered}\mathcal{A}_{i}(\bm{X}(z))=\begin{cases}\kappa^{\mathcal{A}_{i}}_{1}z+\rho^{\mathcal{A}_{i}}_{1}&\text{if }[L^{\mathcal{A}_{i}}_{1},U^{\mathcal{A}_{i}}_{1}],\\ \kappa^{\mathcal{A}_{i}}_{2}z+\rho^{\mathcal{A}_{i}}_{2}&\text{if }[L^{\mathcal{A}_{i}}_{2},U^{\mathcal{A}_{i}}_{2}],\\ \hskip 28.45274pt\vdots\\ \kappa^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}z+\rho^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}&\text{if }[L^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})},U^{\mathcal{A}_{i}}_{K(\mathcal{A}_{i})}],\end{cases}\end{gathered}$ (21)
where $K(\mathcal{A}_{i})$ is the number of linear pieces of the piecewise-linear function, and $\kappa^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ and $\rho^{\mathcal{A}_{i}}_{k}\in\mathbb{R}$ for $k\in[K(\mathcal{A}_{i})]$ are the coefficient and the constant of the $k$-th polyhedron, respectively. For each $i\in[n]$, there exists $k\in[K(\mathcal{A}_{i})]$ such that $z\in[L^{\mathcal{A}_{i}}_{k},U^{\mathcal{A}_{i}}_{k}]$, then the inequality $\mathcal{A}_{i}(\bm{X}(z))\geq\lambda$, can be solved as
$\begin{gathered}\coloneqq\begin{cases}\left[\max\left(L^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right),U^{\mathcal{A}_{i}}_{k}\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}>0,\\ \left[L^{\mathcal{A}_{i}}_{k},\min\left(U^{\mathcal{A}_{i}}_{k},\left((\lambda-\rho^{\mathcal{A}_{i}}_{k})/\kappa^{\mathcal{A}_{i}}_{k}\right)\right)\right]&\text{if }\kappa^{\mathcal{A}_{i}}_{k}<0.\end{cases}\end{gathered}$ (22)
We denote the over-conditioned interval as
$\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)=\bigcap_{i\in[n]}\left[L^{i}_{z},U^{i}_{z}\right].$ (23)
Piecewise Linearity of Diffusion Models.
We now show that the diffusion model mapping $\mathcal{D}$ and then filtered reconstruction error $E$ can be expressed as a piecewise-linear function of $\bm{X}$. To show this, we see that both the forward process and reverse process of the diffusion model are piecewise-linear functions as long as we employ a class of U-Net described below. It is easy to see the piecewise-linearity of the forward process as long as we fix the random seed for $\epsilon_{t}$. To make the reverse process a piecewise-linear function, we employ a U-Net architecture composed of piecewise-linear components such as ReLU activation function and average pooling. Then, $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$ is represented as a piecewise-linear function of $\mathbf{x}_{t}$. Moreover, since $f^{(t)}_{\theta}(\mathbf{x}_{t})$ in (3) is a composite function of $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$, it is also a piecewise-linear function. By combining them together, we see that $\mathbf{x}_{t-1}$ is written as a piecewise-linear function of $\mathbf{x}_{t}$. Therefore, the entire reconstruction process is a piecewise-linear function since it just repeats the above operation multiple times (see Algorithm 1). As a result, the entire mapping ${\mathcal{D}}(\mathbf{X})$ of the diffusion model is a piecewise-linear function of the input image $\bm{X}$. Moreover, since the averaging filter $\mathcal{F}$ and the absolute operation are also piecewise-linear functions, $|\mathcal{F}(\bm{X}-\mathcal{D}(\bm{X}))|(=E(\bm{X}))$ is piecewise-linear. By exploiting this piecewise-linearity, the interval $\mathcal{Z}^{\text{oc}}$ can be computed.
4.3 Identification of $\mathcal{Z}$ by Parametric Programming
Over-conditioning causes a reduction in power due to excessive conditioning. A technique called Parametric Programming is utilized to explore all intervals along the one-dimensional line, resulting in (17). The truncated region $\mathcal{Z}$ can be represented using $\mathcal{Z}^{\mathrm{oc}}$ as
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (24)
The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them. In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z^{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z^{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X})$. An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
Algorithm 2 Selective $p$-value Computation by Parametric Programming
0:  $\mathbf{X^{\mathrm{obs}}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z^{\mathrm{obs}}:=T(\bm{X}^{\mathrm{obs}})$
1:  $\mathcal{Z}\leftarrow\emptyset$
2:  Obtain $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ by (7)
3:  Compute $\bm{a}$, $\bm{b}$ by (17)
4:  $z\leftarrow z_{\mathrm{min}}$
5:  while $z<z_{\mathrm{max}}$ do
6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$ and $\mathcal{M}_{\bm{a}+\bm{b}z}$ by (23) for $z$
7:     if $\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ then
8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$
9:     end if
10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})}+\varepsilon$, where $\varepsilon$ is small positive number.
11:  end while
12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
12:  $p_{\mathrm{selective}}$
5 Experiments
We compared our proposed methods (DMAD-test, DMAD-test-oc) with the other methods: naive method (naive), bonferroni correction (bonferroni), and permutation test (permutation) on type I error rate and power. The details of the methods for comparison are described in Appendix B. We executed the experiment on AMD EPYC 9474F processor, 48-core $3.6$GHz CPU and $768$GB memory.
5.1 Numerical Experiments
Experimental Setup.
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. The details of the structure of the diffusion model we used are shown in Appendix C. The diffusion models were trained on data tailored to each experimental settings. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=200$, and the reconstruction was conducted $5$ step samplings. The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In this experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$. We used only normal images for the type I error rate experiments. The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{0},\Sigma)$. we made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only anomaly images. We made 1000 anomaly images $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{\mu},\Sigma)$, where $\mu_{i}=\Delta,\forall i\in\mathcal{S}$, and $\mu_{i}=0,\forall i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomaly region with its position randomly chosen. The image size of the anomaly images was set to 256, with signals $\Delta\in\left\{1,2,3,4\right\}$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $2$. All experiments were conducted under the significance level $\alpha=0.05$. In addition, we conducted robustness experiments against non-Gaussian noise, i.e., $\bm{X}\not\sim\mathcal{N}$. The details are described in Appendix D.
Results.
Figure 2 shows type I error rate. The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 3 shows powers. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed method has the highest power. DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
(a) Independence
(b) Correlation
Figure 2: Type I Error Rate comparison
(a) Independence
(b) Correlation
Figure 3: Power comparison
5.2 Real Data Experiments
We used the dataset in (Buda et al., 2019). This dataset includes brain images with tumors as normal images (939 images) and brain images without tumors as anomaly images (941 images). A random selection of 100 images from the normal images was made to estimate the mean and variance. The remaining images were standardized using the estimated mean and variance, and 750 normal images were extracted for training the diffusion model. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=330$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.8$ and the kernel size of the averaging filter to $2$. The results of the DMAD-test and naive are shown in the Figure 4. The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors. The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for anomaly images (True Positive), indicating that the results are desirable. On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for anomaly images but also for normal images (False Positive), indicating the invalidness of the naive test. This result indicates that the DMAD-test detected anomaly region as statistically significant while avoiding misidentification of normal image as anomaly.
$p_{\rm naive}=0.027$, $p_{\rm selective}=0.866$ Refer to caption $p_{\rm naive}=0.028$, $p_{\rm selective}=0.611$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.530$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.784$ Refer to caption $p_{\rm naive}=0.001$, $p_{\rm selective}=0.937$ Refer to caption $p_{\rm naive}=0.009$, $p_{\rm selective}=0.145$
(a) Brain images without tumor
$p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.032$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.036$
(b) Brain images with tumor
Figure 4: Example results of the proposed DMAD-test and the naive method applied to brain images.
6 Conclusions, Limitations and Future Works
In this study, we proposed a statistical test for anomaly regions in medical images detected by a diffusion model. With the proposed DMAD-test, the false detection probability can be controlled with the desired significance level because statistical inference is properly conducted conditional on the fact that the anomaly regions are identified by a diffusion model. We demonstrated the proposed method on a simple anomaly detection task, but its applicability to other tasks remains unclear. Moreover, as the complexity of the diffusion model architecture increases, so does the computational cost. Therefore, to apply the proposed method to practical problems, it is necessary to improve computational efficiency.
Acknowledgement
This work was partially supported by MEXT KAKENHI (20H00601), JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
Appendix A Accelerated Reverse Processes
Methods for accelerating the reverse process have been proposed in DDPM, DDIM (Song et al., 2022). When taking a strictly increasing subsequence $\tau$ from $\{1,\cdots,T\}$, it is possible to skip the sampling trajectory from $\mathbf{x}_{\tau_{i}}$ to $\mathbf{x}_{\tau_{i-1}}$. In this case, equations (2) and (4) can be rewritten as
$\mathbf{x}_{\tau_{i-1}}=\sqrt{\alpha_{\tau_{i-1}}}\left(\frac{\mathbf{x}_{\tau_{i}}-\sqrt{1-\alpha_{\tau_{i}}}\cdot\epsilon^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})}{\sqrt{\alpha_{\tau_{i}}}}\right)+\sqrt{1-\alpha_{\tau_{i-1}}-\sigma_{\tau_{i}}^{2}}\cdot\epsilon_{\theta}^{(\tau_{i})}(\mathbf{x}_{\tau_{i}})+\sigma_{\tau_{i}}\epsilon_{\tau_{i}},$ (25)
where
$\sigma_{\tau_{i}}=\eta\sqrt{(1-\alpha_{\tau_{i-1}})/(1-\alpha_{\tau_{i}})}\sqrt{1-\alpha_{\tau_{i}}/\alpha_{\tau_{i-1}}}.$ (26)
Therefore, piecewise-linearity is preserved, making the proposed method DMAD-test applicable.
Appendix B Comparison Methods for Numerical Experiments
We compared our proposed method with the following methods:
  • DMAD-test: The proposed method uses the parametric programming.
  • DMAD-test-oc: The proposed method uses over-conditioning.
  • naive: The naive method. This method uses a conventional $z$-test without any conditioning. The naive $p$-value is calculated as
    $p_{\mathrm{naive}}=\mathbb{P}_{\mathbb{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\right).$
  • bonferroni: To control the type I error rate, this method applies the bonferroni correction. Given that the total number of anomaly regions is $2^{n}$, the $p$-value is calculated as ,
    $p_{\mathrm{bonferroni}}=\min(1,2^{n}\cdot p_{\mathrm{naive}}).$
  • permutation: This method uses a permutation test with the steps outlined below:
    • Calculate the observed test statistic $z^{\mathrm{obs}}$ by applying the observated image $\bm{X}^{\mathrm{obs}}$ to the diffusion model.
    • For each $i=1,\ldots,B$, compute the test statistic $z^{(i)}$ by applying the permuted image $\bm{X}^{(i)}$ to the diffusion model, where $B$ represents the total number of permutations, set to 1,000 in our experiments.
    • The permutation $p$-value is then determined as
      $p_{\mathrm{permutation}}=\frac{1}{B}\sum_{b\in[B]}\bm{1}\{|z^{(b)}|>|z^{\mathrm{obs}}|\},$
      where $\bm{1}\{\cdot\}$ denotes the indicator function.
This rephrasing aims to maintain the original meaning while enhancing readability and comprehension.
Appendix C Structure of the U-Net
Figure 5 shows the architecture of the U-Net used in our experiments. The U-Net has three skip connections, and the Encoder and Decoder blocks. The U-Net employs a dense layer for embedding time $t$.
Figure 5: The architecture of the U-Net
Appendix D Robustness Experiments
To evaluate the robustness of our proposed method against non-Gaussian noise, we conducted experiments using various non-Gaussian distributions with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$.
D.1 Experimental Setup
We considered the following non-Gaussian distributions, each with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • skew-norm: Skew normal distribution.
  • exponnorm: Exponentially modified normal distribution.
  • gennormsteep: Generalized normal distribution.(with a steepness parameter $\beta$ to be steeper than the normal distribution, i.e., $\beta<2$).
  • gennormflat: Generalized normal distribution (with a fatness parameter $\beta$ to be fatter than the normal distribution, i.e., $\beta>2$).
  • t: Student’s $t$-distribution.
Note that these distributions are standardized in the experiments. Figure 6 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. We conducted experiments for each distribution family and each 1-Wasserstein distance $1000$ times to calculate the type I error rate. The significance levels $\alpha$ were set to $0.05$ and $0.10$.
Figure 6: Non-Gaussian distributions with $d=0.04$
D.2 Results
The results of the robustness experiments are shown in Figure 7. Our proposed method, DMAD-test, maintains good performance on type I error rate for all the considered distributions.
(a) $\alpha=0.05$
(b) $\alpha=0.10$
Figure 7: Type I error rate for non-Gaussian distributions